### v5

@@ 7 Discussions
The methodologies employed, grounded in network and game theories, focus on contemporaneously interactive relationships among economies and are devoid of parameters and pre-set utility functions. These network game approaches can be applied to other bilateral flow or dyadic data (such as traffic, social networks, and global payment systems) and similar conflicts (such as currency wars and technology wars).
@@
Several critical factors warrant further investigation. First, counterbalance disequilibrium may arise when trades are notably imbalanced. This imbalance can manifest in various ways. For instance, when country $i$ uses borrowed money to consume country $j$’s exports, country $i$’s consumption still exerts influence over country $j$’s production. However, this influence originates not only from country $i$, but also from the money lender, who possesses some control over $i$’s consumption. If the lender happens to be country $j$, then $\pi_{i}$ is overestimated and $\pi_{j}$ is underestimated, potentially leading to a debt trap for country $i$.
@@
Secondly, unlike an economic superpower, an emerging economy may prioritize the growth of exports over competitive advantages. Indeed, its economic growth must significantly surpass that of advanced economies to bridge the competitiveness gap, in accordance with $\pi$’s Matthew effect. Thus, maximizing competitive advantages may be a low priority for an emerging economy. For example, during China’s primitive capital accumulation in the 1980s and 1990s, earning additional US dollars was essential to finance its industrial modernization. The time dimension, absent from the current research, is necessary to formulate growth strategies.
@@
Thirdly, the facets of economic globalization encompassing capital, information, and technology have yet to be fully integrated. These elements interact with goods and services, but their data are not as easily accessible. Additionally, several data issues may undermine the empirical study’s results. Notably, there are numerous missing values in the service sectors, and discrepancies often arise between country $i$’s reported imports from country $j$ and country $j$’s reported exports to country $i$.
@@
Furthermore, the integration of regional economies warrants further examination. Theorem 2 could be employed to identify suitable candidates for integration. However, the underlying assumption may need modification. For instance, if country $i$ is a member of multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries as used in Theorem 6.
@@
Moreover, additional variables could be incorporated into the demand-sided counterbalance Eq. (3), particularly those from the supply side, such as labor mitigation, which may fine-tune the fractions in $P$. For example, incorporating labor loss into $g_{i}$. Any factors directly influencing country $i$’s production could be included in the $i$th row of $P$. Although consumption may still predominantly drive the fractions, not all production is driven by consumption.
@@
The framework can be extended in several directions. Firstly, amid a bilateral trade war and anti-globalization sentiments, a country may target specific groups of partners without directly impacting others. For instance, as illustrated in Table II, the USA could simultaneously increase tariffs on imports from China, Germany, and Japan.
@@
Secondly, while the dynamical system described by Eqs. (9)–(11) or Eqs. (17)–(18) assumes a fixed $P$, introducing a temporal dimension to $P$ could provide valuable insights. Incorporating a time horizon would facilitate the examination of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and the improvement of trade deficits. Additionally, the temporal dimension could enhance the model’s forecast capabilities. Intertemporal delays often occur when production takes place in one year but consumption in another. If production spans two years, the previous year’s imports of intermediate goods could be included in the current year’s completed production.
@@
Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios. An industrial or geopolitical analysis could provide valuable insights into this choice. For instance, as Japan and South Korea competed in automobile and electronic exports in 2000, $\pi_{j}/\pi_{i}$ was appropriate for $\lambda_{ji}$. In the same year, no industry served as a battleground for China and the USA, making $g_{j}/g_{i}$ a natural choice for $\lambda_{ji}$. Moreover, the literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the importance of the distance between trade partners and the type of merchandise as crucial determinants of trade flows. Consequently, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, partially removing location advantage from $\pi$.
@@
Econometric analyses could also provide additional insights. In the linear regression of Eq. (8), for example, if the estimated residual $\hat{\epsilon}_{i}$ is greater or less than 00, we may infer that $\pi_{i}$ is over- or under-valued, respectively. Including additional regressors, such as debt to GDP ratio, in the regression helps to better understand $\pi$ through explanatory data.
@@
Finally, anticipating the side effects of country $i$’s trade war against country $j$, country $j$ might seek cooperation with third parties that would substantially suffer from the conflict. However, adding more countries to both sides could result in a global trade war, akin to the Cold War between 1947 and 1991.
@@
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$. Accordingly, $P$ has three other simultaneous changes: $-\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $-\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; $-1$ at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $-\lambda_{ji}$ at $(i,i)$. Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
@@
Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the $(i+1)$st row, making the $(i+1)$st row of Eq. (A.5) a zero vector. After dropping the $(i+1)$st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ 0&I_{n-1}\end{array}\right]$ (A.7)
@@
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
@@
Similarly, after we add all rows (except for the first one) in Eq. (A.5) to the $(j+1)$st row, the $(j+1)$st row becomes a zero vector. After dropping the $(j+1)$st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ 0&I_{n-1}\end{array}\right]$
@@
+++++++ + + +++++ + $+$ zero net --
$+$ and -- for net trade surplus and deficit, respectively.
Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, $+$, or --.
@@
Country $j$’s direct competition with $i$ for a higher $\pi_{j}$ results in $\lambda_{ji}=\pi_{j}/\pi_{i}$. If it seeks a better trade balance instead, then the Nash solution expects $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is likely larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Consequently, by Eq. (10), $\pi_{j}$ rises as $j$’s exports to $i$ grow, regardless of trade surplus or deficit.
Similarly, if country $j$ competes with $i$ for a better trade balance, then the solution $\lambda_{ji}=g_{j}/g_{i}>\pi_{j}/\pi_{i}$ and a decreasing $P_{ji}$ leads to a declining $\pi_{j}$ in Eq. (10). Thus, $j$ should also not compete with $i$.
@@
As the last column of Eq. (A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the $(i+1)$st row to make the $(i+1)$st row a zero vector. After dropping the $(i+1)$st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
@@
$-\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
@@
is a zero vector except for its $i$th element. Thus, for any $t\neq i$, the zero value in the $t$th element implies that:
$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}P_{it}.$
@@
$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}=\pi_{i}\sum_{t\neq i}P_{it},$
@@
$\sum\limits_{j\neq i}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}-\pi_{i}P_{ii}$ (A.21)
and:
$\sum\limits_{j=1}^{n}\pi_{j}\left[\Lambda\odot P\right]_{ji}^{\top}=\pi_{i}.$

### v6

@@ 7.2 Discussions and Extensions
The methodologies in this study — rooted in network and game theory — emphasize real-time interactions among economies without relying on fixed parameters or predefined utility functions. These network game approaches can also be applied to other bilateral or dyadic systems, such as traffic networks, social platforms, and global payment infrastructures, as well as analogous conflicts, including currency and technology wars.
@@
Several critical issues warrant further exploration. First, significant trade imbalances can create counterbalance disequilibrium. For example, when country $i$ uses borrowed funds to purchase exports from country $j$, $i$’s consumption still influence $j$’s production. However, this influence originates not only from $i$ but also from the lender, who exerts some control over $i$’s consumption. If the lender is country $j$ itself, then $\pi_{i}$ may be overestimated and $\pi_{j}$ underestimated, potentially leading to a debt trap for $i$.
@@
Second, unlike economic superpowers, emerging economies may prioritize export growth over competitive advantages. To close the competitiveness gap, their economic growth must significantly outpace that of advanced economies, consistent with the Matthew effect in $\pi$. Consequently, maximizing competitive advantages may be a low priority. For instance, during China’s early capital accumulation in the 1980s and 1990s, earning additional U.S. dollars was crucial to finance industrial modernization, such as high-speed rail technology transfer. Incorporating the time dimension — currently absent from this research — is essential for developing effective growth strategies.
@@
Thirdly, key dimensions of economic globalization — namely capital, information, and technology — remain insufficiently integrated into the analysis. Although these elements interact with goods and services, their associated data are comparatively less accessible. Additionally, several data limitations may hinder empirical assessments, including missing values in service sectors and frequent discrepancies between country $i$’s reported imports from country $j$ and $j$’s reported exports to $i$.
@@
The integration of regional economies also warrants deeper investigation. Theorem 2 and $\mathcal{N}_{i}^{+}$ could be employed to identify suitable candidates for integration; however, the underlying assumptions may require revision. For example, if country $i$ participates in multiple RTAs, then $\Delta P_{ii}$ may not be proportionally offset by other countries, as assumed in Theorem 6.
@@
Moreover, additional variables could be incorporated into the demand-side counterbalance, particularly from the supply side — such as labor mitigation — which could refine the fractions in $P$. For instance, accounting for labor loss in $g_{i}$ would capture factors directly influencing country $i$’s production. These factors could be reflected in the $i$th row of $P$. While consumption may still predominantly drive these fractions, not all production is consumption-driven.
@@
The framework offers several avenues for extension. First, in the context of a bilateral trade war and rising anti-globalization sentiments, a country may target specific groups of partners without directly affecting others. For example, as shown in Table II, the USA could simultaneously raise tariffs on imports from China, Germany, and Japan.
@@
Second, while the dynamical systems described by Eqs. (9)–(11) or (17)–(18) assume a fixed $P$, introducing a temporal dimension could yield valuable insights. Incorporating a time horizon would enable analysis of a trade war’s effects on inflation, $\pi$’s dependence on historical values, and improvements in trade deficits. This temporal extension could also enhance the model’s forecasting capabilities. Intertemporal delays often arise when production and consumption are distributed across multiple years. For instance, if production spans two years, imports of intermediate goods from the previous year could be included in the current year’s completed production.
@@
Next, we could allow $\lambda_{ji}$ to depend on additional determinants beyond competitiveness and production ratios. Industrial or geopolitical analyses may guide this choice. For instance, in 2000, Japan and South Korea were competitors in automobile and electronic exports, making $\pi_{j}/\pi_{i}$ a suitable basis for $\lambda_{ji}$. In contrast, no major industry served as a battleground between China and the USA that year, suggesting that $g_{j}/g_{i}$ or $\lambda_{ji}^{*}$ might be more appropriate.
The literature (e.g., Allen, Arkolakis, and Takahashi, 2020; Isard, 1954; Yilmazkuday, 2021) highlights the significance of distance between trade partners and the type of merchandise as key determinants of trade flows. Accordingly, one might discount the $i$th row of $P$ by the distances between country $i$ and its trade partners, thereby partially removing locational advantages from $\pi$.
@@
Econometric analyses could also offer further insights. For example, in the linear regression of Eq. (8), a positive or negative estimated residual $\hat{\epsilon}_{i}$ may indicate that $\pi_{i}$ is overvalued or undervalued, respectively. Including additional regressors, such as the debt-to-GDP ratio, could enhance the explanatory power of $\pi$.
@@
Finally, in anticipation of the side effects of a trade war between country $i$ and country $j$, the latter might seek cooperation with third parties that would also be adversely affected. Country $i$ could also form an alliance with those who would benefit from the conflict. However, expanding the number of countries involved on both sides could escalate the situation into a global trade war, reminiscent of the Cold War era (1947–1991).
@@
We let $P_{ji}$ have a small change $\Delta P_{ji}$ and attempt to calculate its effect on $\pi$, including $\pi_{i}$, $\pi_{-i}$, and $\pi_{j}$. Accordingly, $P$ has three other simultaneous changes: $\Delta P_{ji}$ on $P_{jj}$ to offset the change on $P_{ji}$ in the $j$th row; $\lambda_{ji}\Delta P_{ji}$ on $P_{ij}$ for country $j$’s retaliation upon $i$’s change at $P_{ji}$; and $\lambda_{ji}\Delta P_{ji}$ on $P_{ii}$ to maintain the unit sum of the $i$th row. Therefore, $\frac{\mathrm{d}P}{\mathrm{d}P_{ji}}=\lim\limits_{\Delta P_{ji}\to 0}\frac{\Delta P}{\Delta P_{ji}}$ is a zero $n\times n$ matrix except for: $1$ at $(j,i)$; 1-1 at $(j,j)$; $\lambda_{ji}$ at $(i,j)$; and $\lambda_{ji}$ at $(i,i)$. Without loss of generality, we assume $1\leq i<j\leq n$. Dividing Eq. (A.2) by $\Delta P_{ji}$ and letting $\Delta P_{ji}\to 0$, we get the following equation for the derivative of $\pi$ with respect to $P_{ji}$:
@@
Since all rows of Eq (A.5), except for the first one, sum up to a zero vector, we add them to the (i+1)(i+1)st row, making the (i+1)(i+1)st row of Eq. (A.5) a zero vector. After dropping the (i+1)(i+1)st row in Eq. (A.5) and moving the $i$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{i})^{-1}\\ \vec{0}_{n-1}&I_{n-1}\end{array}\right]$ (A.7)
@@
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ji}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ji}}=(\lambda_{ji}\pi_{i}-\pi_{j})\gamma_{ji}$
@@
Similarly, after we add all rows (except for the first one) in Eq. (A.5) to the (j+1)(j+1)st row, the (j+1)(j+1)st row becomes a zero vector. After dropping the (j+1)(j+1)st row in Eq. (A.5) and moving the $j$th column to the first column without changing the order of other columns, we obtain the augmented matrix for $\left(\frac{\mathrm{d}\pi_{j}}{\mathrm{d}P_{ji}},\frac{\mathrm{d}\pi_{-j}^{\top}}{\mathrm{d}P_{ji}}\right)^{\top}$:
@@
$\left[\begin{array}[]{cc}1&-\vec{1}_{n-1}^{\top}(I_{n-1}-Z_{j})^{-1}\\ \vec{0}_{n-1}&I_{n-1}\end{array}\right]$
@@
++++++ ++++ ++ zero net -
++ and - for net trade surplus and deficit, respectively.
Magnitudes in the same row are indicated by the numbers of $\downarrow$, $\uparrow$, ++, or -.
@@
When country $j$ directly competes with country $i$ to increase $\pi_{j}$, the Nash solution yields $\lambda_{ji}=\pi_{j}/\pi_{i}$. In this case, $\pi_{j}$ remains unchanged. However, if country $j$ strategically aims to improve its trade balance instead, the Nash solution becomes $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is typically larger than $\pi_{j}/\pi_{i}$ according to the Matthew effect. Therefore, by Eq. (10), country $j$ achieves its original objective of enhancing $\pi_{j}$, as its exports to $i$ increase — regardless of a trade surplus or deficit.
Similarly, if country $j$ competes with $i$ solely to improve its trade balance, the Nash solution is $\lambda_{ji}=g_{j}/g_{i}$, which results in zero trade surplus. But if country $j$ strategically seeks to enhance $\pi_{j}$, the Nash solution again becomes $\lambda_{ji}=\sqrt{\pi_{j}g_{j}}/\sqrt{\pi_{i}g_{i}}$, which is generally smaller than $g_{j}/g_{i}$ due to the Matthew effect. Consequently, an increasing in $P_{ji}$ leads to a positive trade surplus, fulfilling country $j$’s original objective.
@@
As the last column of Eq. (A.17) sums to zero, i.e., $\sum\limits_{j=1}^{n}\pi\xi_{j}=\pi\sum\limits_{j=1}^{n}\xi_{j}=0$, we add all rows (except for the first) to the (i+1)(i+1)st row to make the (i+1)(i+1)st row a zero vector. After dropping the (i+1)(i+1)st row and moving the $i$th column to the first column without changing the order of other columns, we get the augmented matrix for $\left(\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}},\frac{\mathrm{d}\pi_{-i}^{\top}}{\mathrm{d}P_{ii}}\right)^{\top}$:
@@
$\frac{\mathrm{d}\pi_{i}}{\mathrm{d}P_{ii}}\alpha_{i}+\left(I_{n-1}-Z_{i}\right)\frac{\mathrm{d}\pi_{{}_{-i}}}{\mathrm{d}P_{ii}}=(\pi M_{i})^{\top}$
@@
is a zero vector except for its $i$th element.
For any $t\neq i$, zero value of the $t$th element of the above vector implies that:
$\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\frac{\pi_{i}(1-P_{ii})P_{it}}{1-P_{ii}}=0.$
@@
$\sum_{t\neq i}\sum\limits_{j\neq i}\frac{\lambda_{ij}P_{ij}P_{jt}}{1-P_{ji}}\pi_{j}-\pi_{i}\sum_{t\neq i}P_{it}=0$
@@
$\sum\limits_{j\in\mathcal{N}}\lambda_{ij}P_{ij}\pi_{j}=\pi_{i}.$ (A.21)